# Review of Demographic Bias in Face Recognition

### v1

@@ VII Conclusion
In this work, we have systematically explored the issue of demographic bias in FR through different yet interrelated sections: causes, datasets, assessment metrics, and mitigation strategies. The section on causes delves into the underlying factors, categorizing them into aspects such as imbalances in training datasets, variability of skin-tones, algorithmic sensitivities, image quality and covariates, and combined demographic factors. The discussion on datasets presents an overview of existing resources and their demographic distributions, Furthermore, we reviewed state-of-the-art metrics for evaluating bias, ranging from traditional statistical measures to advanced fairness indicators, highlighting differences in demographic performance and outcomes across demographic groups. The mitigation section highlights diverse approaches, including preprocessing, in-processing, and post-processing techniques, offering a comprehensive overview of existing work in this area.
In addition to summarizing existing research, we identified emerging challenges including the fairness implications of lightweight models, the role of synthetic datasets, the complexities of remote identity verification systems, and intersectional bias. These challenges reflect the evolving nature of FR technologies and underscore the need for innovative strategies to ensure equitable and reliable outcomes in real-world applications.

### v2

@@ VII Conclusion
In this work, we have systematically explored the issue of demographic fairness or bias in FR through different yet interrelated sections: causes, datasets, assessment metrics, and mitigation strategies. We discussed key contributing factors such as training data imbalance, algorithmic sensitivities, and image quality variations, as well as the growing recognition of non-demographic attributes. Our analysis of datasets emphasized the importance of demographic diversity and the need for multi-attribute annotations to enable deeper understanding of performance disparities. We also reviewed a range of fairness evaluation metrics and mitigation techniques across the FR pipeline.
An important challenge to be noted here is the trade-off between fairness and accuracy. Since most FR models are already optimized for performance, adding fairness objectives often introduces conflicting constraints. Although some architectures sometimes improve both, these gains frequently rely on increased model capacity, rather than an inherent resolution of the fairness-accuracy trade-off. Recent studies have shown that gender-related bias may be explained by soft-biometric attributes rather than inherent biological differences, suggesting that some disparities are socio-cultural in nature. However, such understanding is still limited for attributes like race or skin-tone, largely due to the absence of datasets annotated with both demographic and non-demographic variables. Creating such datasets is essential but resource-intensive; synthetic data may help address this gap, though its effectiveness remains sub-par of real-world data. Improved metrics are also needed to quantify the impact of specific attributes on performance in a consistent and interpretable way. Additionally, most fairness evaluations are conducted on large-scale models, yet future deployments–especially on mobile and edge devices–will require lightweight models that may be more prone to differential treatment. Ensuring fairness under such constraints is critical for real-world adoption.
Together, these challenges reflect the evolving nature of FR technologies and underscore the need for innovative strategies to ensure equitable and reliable outcomes in real-world applications.
[Dataset URLs] This appendix provides the URLs for the datasets referenced in Section III. These URLs were accessed in March 2025, corresponding to the time of preparation of this woek. It is important to note that many of these datasets are intended solely for non-commercial research purposes. Readers are advised to review the licensing terms prior to use.